-7 - 7x= -70.... is one of the problem's and she is not able to understand how to work the problem. can we get a better answer to how to figure it out.
Answer by checkley75(3666)

-7-7X=-70 ADD 7 TO BOTH SIDES OF THE EQUATION THUS:
(-7+7)-7X=(-70+7)
0-7X=-63 NOW DIVIDE BOTH SIDES BY -7
-7X/-7=-63/-7
X=9 ANSWER.
PROOF
-7-7*9=-70
-7-63=-70
-70=-70